OpenStudy (anonymous):

solve for t: 3t+1= sqaure root of 7t+1

OpenStudy (anonymous):

(3t+1)^2 = 7t+1 9t^2 + 6t + 1 = 7t + 1 9t^2 + 6t = 7t 9t + 6 = 7 t = 1/9

OpenStudy (anonymous):

where did you get 6t?

OpenStudy (anonymous):

Whenever you expand (3t+1)^2, you get 9t^2 + 6t + 1. (3t+1)^2 shouldn't be confused with 3t^2 + 1^2. (3t+1)^2 is equivalent to (3t+1)(3t+1), where you use the distributive property or FOIL if that's the method you learned in class. (Multiplying first, outer, inner, last) so that you don't forget a pair when expanding quantities to exponents.

OpenStudy (anonymous):

oh gotcha! thanks!

OpenStudy (anonymous):

sure, you're welcome.

OpenStudy (anonymous):

hey i got one more for you...